Solve the following stoichiometry grams-grams problems:




1) Using the following equation:




Pb(SO4)2 + 4 LiNO3 




Pb(NO3)4 + 2 Li2SO4



How many grams of lithium nitrate will be needed to make 250 grams of lithium sulfate,



assuming that you have an adequate amount of lead (IV) sulfate to do the reaction?




2) Using the following equation:




Fe2O3 + 3 H2 -------------> 2 Fe + 3 H2O




Calculate how many grams of iron can be made from 16.5 grams of Fe2O3 by the following



equation.

1
Expert's answer
2022-03-13T15:20:35-0400

1)

"m(LiNO_3)=250\ g(Li_2SO_4)\times\frac{1\ mol(Li_2SO_4)}{109.94\ g(Li_2SO_4)}\times\frac{4\ mol(LiNO_3)}{2\ mol(Li_2SO_4)}\times\frac{68.94\ g(LiNO_3)}{1\ mol(LiNO_3)}=314\ g"


Answer: 314 g


2)

"m(Fe)=16.5\ g(Fe_2O_3)\times\frac{1\ mol(Fe_2O_3)}{159.69\ g(Fe_2O_3)}\times\frac{2\ mol(Fe)}{1\ mol(Fe_2O_3)}\times\frac{55.85\ g(Fe)}{1\ mol(Fe)}=11.5\ g"


Answer: 11.5 g